OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation 4x3+4=0

OpenStudy (mertsj):

1. Subtract 4 from both sides 2. Divide both sides by 4 3. Take the 3rd root of both sides.

OpenStudy (anonymous):

whats the answer @Mertsj

OpenStudy (anonymous):

thankkk Uu @Mertsj

zepdrix (zepdrix):

If you need exact solutions, you can apply your sum of cubes formula, \[\Large\bf\sf 4(x^3+1)=0,\qquad\qquad x^3+1=0\]\[\Large\bf\sf (x+1)(x^2-x+1)=0\]then follow it up with the quadratic formula.

zepdrix (zepdrix):

no big woop, just another option :D

OpenStudy (anonymous):

thanks but i have to turn this in tomoorow and i have a bad grade if u could help me with the answers plz @zepdrix

zepdrix (zepdrix):

\[\Large\bf\sf (x+1)(\color{royalblue}{x^2-x+1})=0\]Apply quadratic formula to the blue part.\[\Large\bf\sf \color{royalblue}{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\] Plug stuff in, \[\Large\bf\sf \color{royalblue}{x=\frac{1\pm\sqrt{1-4}}{2}}\]Simplify.